# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v1

@@ 2.5. AI Model Taxonomy
  • Conventional Machine Learning (C-ML) Models: C-ML models primarily use manual feature extraction and data-driven statistical techniques for pattern recognition. Examples include Decision Trees (DT) (myles2004introduction, ), Support Vector Machines (SVM) (hearst1998support, ), Linear Regression (LR) (montgomery2021introduction, ), Naive Bayes (NB) (webb2010naive, ), K-Nearest Neighbors (KNN) (peterson2009k, ), Random Forest (RF) (breiman2001random, ), etc. Artificial neural networks with very few layers such as 3-layer multilayer perceptrons (MLP) (gardner1998artificial, ) are also viewed as C-ML models. It is noteworthy that deep learning is commonly viewed as an important branch of machine learning, thus we use “conventional machine learning” to denote classic machine learning models apart from deep learning models.
  • Conventional Deep Learning (C-DL) Models: Defined as neural networks with deep layers (typically more than 3 layers) or commonly-used deep learning backbones for automatic feature extraction and representation learning from data. Examples include Convolutional Neural Networks (CNN) (gu2018recent, ), Recurrent Neural Networks (RNN) (medsker2001recurrent, ), vanilla long short-term memory (LSTM) (hochreiter1997long, ), ResNet (he2016deep, ), etc.
  • Advanced Deep Learning (A-DL) Models: Incorporates more complex or novel architectures, advanced training techniques, and elements of other AI fields like reinforcement learning into classic deep networks. Examples include graph convolutional networks (GCN) (kipf2016semi, ), Transformers (vaswani2017attention, ), Generative Adversarial Networks (GANs) (goodfellow2014generative, ), etc.
3. Survey Statistics and Analyses
@@
3.1. Study Selection
We adopt the Preferred Reporting Items for Systematic Reviews and Meta-Analysis (PRISMA) (page2021prisma, ) to conduct our survey, and the flow diagram of PRISMA is provided in the appendices. We summarize the literature selection process as follows: During the paper identification stage, we retrieve 840 records from PubMed, 895 records from WoS, and 596 records from Google Scholar. By comparing the collected records across these three databases, a total of 2069 records remain after 262 duplicate records are removed. In the paper screening stage, after screening the titles and abstracts of the remaining 2069 papers according to the inclusion criteria, 243 articles are further selected for the full-text assessment. Finally, after an in-depth review of the full texts of these selected papers, 103 articles are excluded and 140 regular articles are included in this survey for further information extraction and content summarization. The list of excluded papers from the full-text screening stage, along with the reasons for their exclusion, is detailed in our appendices. Moreover, 14 eligible survey papers (kour2022vision, ; khera2022age, ; fraiwan2021computer, ; kour2019computer, ; kour2023sensor, ; di2020gait, ; pardoel2019wearable, ; gupta2023new, ; figueiredo2018automatic, ; vienne2017inertial, ; das2022recent, ; salchow2022emerging, ; loh2021application, ; ayaz2023automated, ) are found in the screening stage. From these survey papers, we identify another 99 potentially-matched articles via screening their titles and abstracts. After the full-text assessment, we include 24 regular articles from those. In summary, we include a total of 164 eligible articles, comprising 140 from the academic databases and 24 from survey papers, in our survey.
3.2. Statistics of Selected Papers
@@
3.2.1. Distribution of Publication Years
@@ 3.2.2. Distribution of Countries
3.2.3. Distribution of ND Types
Fig. 8b shows the numbers of papers focused on different NDs types. Note that the studies containing diagnosis of multiple NDs (19 articles) are also included into each NDs type ($e.g.$, a study simultaneously diagnosing PD, HD, and ALS are counted as one PD/HD/ALS study). The statistics show that PD receives the most research focus among all the NDs, including a total of 143 articles ($87.2\%$). The second most studied ND’s type is HD with 26 papers ($15.9\%$), while there are 22 ($13.4\%$), 6 ($3.7\%$), 5 ($3.0\%$) papers focused on ALS, MS, and AD respectively. Interestingly, although AD is the most populated diseases among all NDs (see Fig. 2), the study of its diagnosis based on gait has yet to receive much attention, possibly because other NDs patients, such as PD patients, may exhibit more evident abnormal gaits that can be exploited to perform more reliable diagnosis. Promisingly, more and more studies (mannini2016machine, ; daliri2012automatic, ; xia2015classification, ; zhao2021multimodal, ; beyrami2020robust, ; xia2016novel, ; patil2019clinical, ; creagh2020smartphone, ) have demonstrated the feasibility of leveraging gait data to identify HD, AD, ALS, and MS, therefore a deeper exploration of these research directions should be conducted.
3.2.4. Distribution of Gait Data Types
@@
Fig. 8e further shows the distribution of sensor types used in 137 sensor-based studies: The force sensors (FS) ($e.g.,$ force/pressure sensitive resistors) and IMUs are the two most widely-used sensor types, occupying $56.2\%$ (77 studies) and $31.4\%$ (43 studies) respectively. For other sensors (17 studies), there are 5 studies combining both FS and IMUs as gait data sources, while 5 studies utilize gait analysis system (GAS) ($e.g.,$ mats, walkways, treadmills). The low-cost and non-invasive motion capture device, Kinect (seifallahi2022alzheimer, ; munoz2022machine, ), is also used as both a sensor and a vision-based device for gait data collection (2 studies). We provide a further discussion on exploiting sensor-based ($e.g.,$ Kinect-based) 3D skeleton data for NDs diagnosis in Sec. 5.2.
3.2.5. Distribution of AI Models
@@
3.2.6. Distribution of Sample Sizes

### v2

@@ 2.5 AI Model Taxonomy
  • Conventional Machine Learning (C-ML) Models: C-ML models primarily use manual feature extraction and data-driven statistical techniques for pattern recognition. Examples include Decision Trees (DT) [85], Support Vector Machines (SVM) [86], Linear Regression (LR) [87], Naive Bayes (NB) [88], K-Nearest Neighbors (KNN) [89], Random Forest (RF) [90], etc. Artificial neural networks with very few layers such as 3-layer multilayer perceptrons (MLP) [91] are also viewed as C-ML models. It is noteworthy that deep learning is commonly viewed as an important branch of machine learning, thus we use “conventional machine learning” to denote classic machine learning models apart from deep learning models.
  • Conventional Deep Learning (C-DL) Models: Defined as neural networks with deep layers (typically more than 3 layers) or commonly-used deep learning backbones for automatic feature extraction and representation learning from data. Examples include Convolutional Neural Networks (CNN) [92], Recurrent Neural Networks (RNN) [93], vanilla long short-term memory (LSTM) [94], ResNet [95], etc.
  • Advanced Deep Learning (A-DL) Models: Incorporates more complex or novel architectures, advanced training techniques, and elements of other AI fields like reinforcement learning into classic deep networks. Examples include graph convolutional networks (GCN) [96], Transformers [97], Generative Adversarial Networks (GANs) [98], etc.
3 Survey Statistics and Analyses
@@
3.1 Study Selection
We adopt the Preferred Reporting Items for Systematic Reviews and Meta-Analysis (PRISMA) [99] to conduct our survey, and the flow diagram of PRISMA is provided in the appendices. We summarize the literature selection process as follows: During the paper identification stage, we retrieve 840 records from PubMed, 895 records from WoS, and 596 records from Google Scholar. By comparing the collected records across these three databases, a total of 2069 records remain after 262 duplicate records are removed. In the paper screening stage, after screening the titles and abstracts of the remaining 2069 papers according to the inclusion criteria, 243 articles are further selected for the full-text assessment. Finally, after an in-depth review of the full texts of these selected papers, 103 articles are excluded and 140 regular articles are included in this survey for further information extraction and content summarization. The list of excluded papers from the full-text screening stage, along with the reasons for their exclusion, is detailed in our appendices. Moreover, 14 eligible survey papers [46, 34, 35, 47, 36, 37, 38, 39, 100, 40, 41, 42, 43, 44] are found in the screening stage. From these survey papers, we identify another 99 potentially-matched articles via screening their titles and abstracts. After the full-text assessment, we include 24 regular articles from those. In summary, we include a total of 164 eligible articles, comprising 140 from the academic databases and 24 from survey papers, in our survey.
3.2 Statistics of Selected Papers
@@
3.2.1 Distribution of Publication Years
@@ 3.2.2 Distribution of Countries
3.2.3 Distribution of ND Types
Fig. 8b shows the numbers of papers focused on different NDs types. Note that the studies containing diagnosis of multiple NDs (19 articles) are also included into each NDs type ($e.g.$, a study simultaneously diagnosing PD, HD, and ALS are counted as one PD/HD/ALS study). The statistics show that PD receives the most research focus among all the NDs, including a total of 143 articles ($87.2\%$). The second most studied ND’s type is HD with 26 papers ($15.9\%$), while there are 22 ($13.4\%$), 6 ($3.7\%$), 5 ($3.0\%$) papers focused on ALS, MS, and AD respectively. Interestingly, although AD is the most populated diseases among all NDs (see Fig. 2), the study of its diagnosis based on gait has yet to receive much attention, possibly because other NDs patients, such as PD patients, may exhibit more evident abnormal gaits that can be exploited to perform more reliable diagnosis. Promisingly, more and more studies [48, 101, 102, 77, 103, 104, 105, 106] have demonstrated the feasibility of leveraging gait data to identify HD, AD, ALS, and MS, therefore a deeper exploration of these research directions should be conducted.
3.2.4 Distribution of Gait Data Types
@@
Fig. 8e further shows the distribution of sensor types used in 137 sensor-based studies: The force sensors (FS) ($e.g.,$ force/pressure sensitive resistors) and IMUs are the two most widely-used sensor types, occupying $56.2\%$ (77 studies) and $31.4\%$ (43 studies) respectively. For other sensors (17 studies), there are 5 studies combining both FS and IMUs as gait data sources, while 5 studies utilize gait analysis system (GAS) ($e.g.,$ mats, walkways, treadmills). The low-cost and non-invasive motion capture device, Kinect [107, 82], is also used as both a sensor and a vision-based device for gait data collection (2 studies). We provide a further discussion on exploiting sensor-based ($e.g.,$ Kinect-based) 3D skeleton data for NDs diagnosis in Sec. 5.5.
3.2.5 Distribution of AI Models
@@
3.2.6 Distribution of Sample Sizes
@@
4 Overview of Existing Studies
In this section, we systematically review existing studies related to AI-assisted NDs diagnosis based on human gait. First, a quality evaluation criterion based on the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples, is devised to quantitatively measure the quality of each study. Then, we categorize all studies by different NDs types (PD, HD, AD, ALS, MS) with a data summary for them (shown in Table 2, 3, 4, 5, 6, and 7). Last, according to the quality scores, we provide a discussion on representative studies.
4.1 Quality Evaluation Criterion
To comprehensively measure the quality of each study, we empirically grade the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples at three levels with the score 1, 2, or 3. We sum their scores and divide them by the maximum score to get a normalized quality score $S_{Q}$ ranging from 0 to 1. $S_{Q}$ is computed for each study, which are defined as follows:
  • Score ($S_{n}$) for Novelty of AI Models
    1. 1.
      Score $S_{n}=1$: The main proposed method belongs to or contains only C-ML models.
    2. 2.
      Score $S_{n}=2$: The main proposed method belongs to or contains C-DL models.
    3. 3.
      Score $S_{n}=3$: The main proposed method belongs to or contains A-DL models.
  • Score ($S_{c}$) for Comprehensiveness of Method Comparison
    1. 1.
      Score $S_{c}=1$: The study only compares different settings of the proposed method, without comparing with any other existing methods or studies.
    2. 2.
      Score $S_{c}=2$: The study compares the proposed method with one published method or study.
    3. 3.
      Score $S_{c}=3$: The study compares the proposed method with more than one published methods or studies.
  • Score ($S_{s}$) for Sufficiency of Experimental Samples
    1. 1.
      Score $S_{s}=1$: The size of experimental samples ($i.e.$, number of patients plus number of controls) is less than 30.
    2. 2.
      Score $S_{s}=2$: The size of experimental samples is equal to or more than 30, and less than 100.
    3. 3.
      Score $S_{s}=3$: The size of experimental samples is equal to or more than 100.
To normalize the final quality score $S_{Q}$ to the value range from 0 to 1, we sum the scores from all three aspects and divide them by the maximum score:
$S_{Q}=\frac{S_{n}+S_{c}+S_{s}}{9}$ (1)
The quality score $S_{Q}$ is computed for each study and presented in the tables. We empirically set $S_{Q}>0.6$ as the threshold for selecting representative high-quality studies.
Table 1: Number of studies related to different NDs (PD, HD, MS, AD, ALS) and different model types (C-ML, C-DL, A-DL). “Combinatorial” denotes studies that simultaneously contain diagnosis of multiple NDs including PD, ALS, and HD. $S_{Q}>0.6$ indicates the number of high-quality studies.
        Type         C-ML         C-DL         A-DL         Total         $\boldsymbol{S_{Q}>0.6}$
        PD         77         34         13         124         56
        Combinatorial
        (PD, ALS, HD)
        11         6         2         19         8
        HD         6         1         0         7         1
        MS         4         1         1         6         2
        AD         3         1         1         5         2
        ALS         3         0         0         3         0
        Total         104         43         17         164         —
        $\boldsymbol{S_{Q}>0.6}$         23         29         17         —         69
4.2 Parkinson’s Disease (PD)
As shown in Table 2, 3, 4, 5, and 6, there are 124 studies that focus on PD diagnosis based on human gait. It is worth mentioning that there are 19 studies, each of which simultaneously includes the diagnosis of PD and other NDs, as shown in Table 6 (referred to as “Combinatorial Diagnosis of Different NDs”). We separately discuss these studies in Sec. 4.3. Based on our survey, there are 77, 34, and 13 PD studies using C-ML, C-DL, and A-DL models respectively. It should be noted that there are a large number of PD studies (up to 124 studies). Since the studies with high-quality AI models and diagnostic results are the focus of our survey, we first categorize all studies by three AI model types (C-ML, C-DL, A-DL), and then select representative studies with quality score $S_{Q}>0.6$ for content summary and analysis.
4.2.1 C-ML Models for PD Diagnosis
[108] (Table 2, ID = P-7) utilizes Hidden Markov Models (HMM) with Gaussian Mixtures to classify gait data and differentiate Parkinson’s disease (PD) patients from healthy subjects. Using stride interval as the distinguishing feature, the HMM method achieves an accuracy rate of 90.3%. In [80] (ID = P-10), the authors presents a methodology using Microsoft Kinect sensors to detect gait attributes and recognize gait of PD patients. The leg length, stride length, and gait velocity are extracted to achieve high-accuracy classification using neural networks (NN). This study suggests the potential of using Kinect sensors for gait analysis and disorder recognition. [109] (ID = P-14) proposes a radial basis function neural networks (RBF-NN) based method for classifying PD patients and healthy individuals using gait characteristics derived from vertical ground reaction forces (vGRF), achieving over 90% accuracy. [110] (ID = P-16) focuses on a technique called Shifted One-Dimensional Local Binary Patterns (Shifted 1D-LBP) to track local changes in time series data. Combining with machine learning model, this work demonstrate Shifted 1D-LBP’s potential in real-time NDD detection applications. The study of [57] (ID = P-20) utilizes the vGRF data to extract gait features and feeds them into the SVM classifier to obtain a good accuracy result of 93.6%. [111] (ID = P-21) explores the use of wavelet analysis combined with SVM for the identification of PD based on spatio-temporal gait variables, achieving a classification accuracy of 90.32%. The study demonstrates the potential of using wavelet analysis as an efficient method for classifying PD and healthy subjects. [112] (ID = P-30) utilizes a wearable inertial device to collect motor data from healthy subjects, individuals with idiopathic hyposmia (IH), and PD patients during lower limb assessments, and develops a machine learning-based system with SVM, RF, NB to classify the three groups. In [113] (ID = P-32), the authors quantify gait parameters objectively using wearable accelerometers, and compare them with a motion capture system to automatically discriminate patients with PD using machine learning models such as SVM, KNN, and DT. The proposed approach achieves an accuracy of approximately 89.1% and suggests that the wearable accelerometer-based system is suitable for assessing and monitoring PD and Freezing of gait (FOG) in real-life scenarios. [114] (ID = P-38) emphasizes of using a hybrid of linear and non-linear features extracted from human gait time series data to improve the NDD classification system which employs SVM with a radial basis kernel function. [115] (ID = P-42) leverages different classic machine learning methods (KNN, DT, RF, NB, SVM, K-Means) to identify PD from vGRFs collected from gait cycles. The structure of PD recognition process from data acquisition to performance evaluation is also systematically presented in this study. [116] (ID = P-47) devises an automated system with force sensors to extract gait parameters and then use these parameters as input to the SVM model to diagnosis PD. [117] (ID = P-50) proposes a novel method for classifying gait patterns between patients with PD and healthy controls. This method reconstructs the phase space of vGRF and decomposes the gait dynamics using empirical mode decomposition, which are fed to neural networks (NN) for classification with accuracy of 98.8%. In [118] (ID = P-57), the authors explore the effectiveness of Extreme Gradient Boosting (XGBoost) and Artificial Neural Network (ANN) models for monitoring the gait of Parkinson’s Disease patients, achieving 41.0%(double check the figure) accuracy. The study of [119] (ID = P-60) explores the use of single-sensor accelerometer data from healthy subjects to extract features relevant to multi-sensor accelerometer gait data for Parkinson’s disease classification. They use a pre-trained convolutional autoencoder as the source model and a simple multi-layer perceptron as the target model to classify PD and achieve 68.92% accuracy. A machine learning based gait classification system with different models such as SVM and RT is presented in [120] (ID = P-65) to assist in diagnosing PD form vGRF. [121] (ID = P-71) extracts spatio-temporal gait features such as stride time and stance time from vGRF and exploits different classic machine learning models (KNN, NB, SVM) to classify health control and PD subjects with an accuracy of 88.5-98.8%. In [122] (ID = P-81), a time-varying singular value decomposition method is designed to extract the most useful and informative part of the vGRF signal data, and then sparse non-negative least squares (NNLS) approach is leveraged to classify PD and HCs, achieving 97.2% accuracy. [123] (ID = P-87) utilizes RF to identify PD patients from vGRF and achieves up to 98.8% accuracy. The SVM model is used in [124] (ID = P-111) to classify PD and HC gait from vGRF data, with an accuracy of 98.2%.
4.2.2 C-DL Models for PD Diagnosis
[125] (ID = P-37) proposes a Locally Weighted Random Forest (LWRF) regression model to predict the PD symptoms using time- and frequency-domain features derived from vGRF signals collected using foot worn sensors. [126] (ID = P-39) develop a hybrid model combining Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN) for PD’s detection and severity rating. The model captures the spatio-temporal characteristics of gait dynamics using vGRF signals and offers a multi-category classification approach for PD severity, beyond the binary detection typically provided by other ML methods. The study of [81] (ID = P-43) leverages convolutional LSTM to identify PD patients from Kinect-based skeleton data and achieves up to 83% accuracy. It also suggests that the data preprocessing ($e.g.$, cropped data) can help boost the classification performance. [127] (ID = P-52) devises a deep neural network (DNN) based model to identify Parkinsonian gait based on vGRF data and achieves 98.7% accuracy. This is also the first study to predict the Unified Parkinson’s Disease Rating Scale (UPDRS) severity ($i.e.$, Parkinsion severity rate) of a subject and achieve 85.3% accuracy. [128] (ID = P-58) utilizes both CNN and attention-enhanced LSTM networks for left and right gait analysis, which helps achieve high-competitive PD classification accuracy (98.03%). [129] (ID = P-63) uses balance and gait variables obtained with wearable inertial motion sensors to differentiate PD patients using machine learning models (NN, SVM, KNN, DT, RF). [130] (ID = P-64) applies a mask R-CNN model for extracting human silhouettes from video frames, and then obtains gait energy images (GEIs) from human silhouettes as features for PD gait classification. [131] (ID = P-66) presents a novel Bidirectional Long Short-Term Memory (Bi-LSTM) model to classify PD from HC using motion data, acquired during performing 13 tasks derived from the Movement Disorder Society-Unified Parkinson’s Disease Rating Scale III (MDS-UPDRS III), with wearable sensors. [132] (ID = P-67) reports a smart-phone based PD diagnosis method by utilizing real-world accelerometer records and achieved an AUC score of 0.86. This method demonstrate the superior accuracy for at-home screening of PD and other NDD conditions affecting movement while acknowledges limitations such as potential over-enrollment of tremor-dominant PD subjects in their experiment. [133] (ID = P-68) uses a novel technique called Neighborhood Representation Local Binary Pattern (NR-LBP) and ANN to model the relationship between statistical features extracted from vGRF data and PD symptoms. [134] (ID = P-69) utilizes LSTM to differentiate the PD patients and healthy subjects based on the vGRF and gait time series, which performs better than other state-of-the-art methods with binary-classification accuracy of 98.6%. The Bi-LSTM model is leveraged in [135] (ID = P-70) to classify PD gait using vGRF signals from single sensors. This study reveals that combining time-frequency and time-space features of physiological signals are more effective and can boost the classification accuracy of deep-learning models. In [136] (ID = P-74), force domain, peak domain, and abnormality domain features (22 features) are combined to identify PD patients based on CNN and XGBoost, with an accuracy of 95.68%-97.32%. [84] (ID = P-76) characterizes the pre-diagnostic period of PD and develops a machine learning model (LR) and deep learning model (RNN) for PD diagnosis. In [137] (ID = P-84), the authors focus on identifying changes in gait patterns associated with PD by analyzing the spatio-temporal data of vGRF. They propose a model based on CNN that evaluates the changes in vGRF and identify the Parkinsonian gait from healthy gait with an accuracy of 97.0%. [63] (ID = P-86) combines CNN and gated recurrent units (GRUs) to perform classification between PD and health subjects from IMU signals, achieving 83.7-92.7% accuracy. The results suggest that the combination of temporal and spectral information is more effective in classifying gait patterns of PD patients. [138] (ID = P-104) devises an anomaly detection algorithm to classify PD and healthy controls based on wrist-sensor recordings (tri-axial accelerometer data) and achieves 83.3% accuracy. [139] (ID = P-115) suggests a real-time PD identification method using wavelet transform-based features extracted from vGRF data. Experimental results using NN with resilient back-propagation algorithm achieved an average efficiency of 86.75%. 2D CNN is leveraged in [140] (ID = P-116) to diagnose PD at an early stage from vGRF signal data and achieves 98% accuracy. In [141] (ID = P-120), a plantar pressure analysis method is proposed to identify the gait dynamics of PD patients using a constant radial basis function neural network (RBFNN).
4.2.3 A-DL Models for PD Diagnosis
In [142] (ID = P-15), a novel hybrid model based on BP neural networks is proposed to classify PD and HC, which outperforms four baseline methods on a public dataset with an average accuracy of 87.9%. [143] (ID = P-17) introduces a Q-backpropagated Time Delay Neural Network (Q-BTDNN), which harvests the temporal gait data collected from wearable devices, combined with the Q-backpropagation (Q-BP) algorithm for predicting gait disturbances in PD. The study of [144] (ID = P-31) proposes a computational method based on tensor decomposition for analyzing gait dynamics in PD from multisensor time series data, achieving an accuracy of 86.2%. [145] (ID = P-75) uses deep CNN to classify gait deterioration in PD patients based on spatio-temporal vGRF signals and achieves 95.5% accuracy in average. The study suggests that the heel strike and body balance are the most indicative gait elements in PD classification. [146] (ID = P-82) develops an NN model to classify PD and HC based on the vGRF dataset collected from a public dataset and achieves an accuracy result of 84.8%. [147] (ID = P-85) proposes an interval type-2 fuzzy neural network with novel quasi-Levenberg-Marquardt (qLM) learning approach to identify PD patients from the gait cycle based vGRF signals. Based on the 10 clinical features extracted from 14 gait features, the model achieves 97.6% accuracy with a smaller size of parameters and a good interpretability ($i.e.$, rules based on fuzzy logic). In [148] (ID = P-93), the authors propose a novel architecture PD-ResNet based on ResNet [95] and use it to differentiate PD and HC. Its result outperforms many existing state-of-the-art models with 95.5% accuracy. A hybrid deep learning model combining CNN and locally weighted RF (LWRF) [149] (ID = P-97) is devised to perform PD detection tasks from time and frequency features of vGRF, achieving 99.5% accuracy. [150] (ID = P-98) presents a Riemannian deep learning strategy called SPDnet to learn compact descriptors for discriminating Parkinsonian gait patterns. The proposed SPDnet learns discriminatory patterns between Parkinson’s and control populations, outperforming traditional 3D convolutional representations with 99.4% accuracy. [77] (ID = P-106) proposes a correlative memory neural network (CorrMNN) scheme to extract spatio-temporal features to discriminate NDDs’ gait from normal gait based on vGRF, and achieves up to 100% accuracy. A video-based SlowFast GCN network is proposed in [151] (ID = P-109) to leverage 2D pose body-joint graph for diagnosing gait disorders such as PD. [152] (ID = P-121) proposes a novel model named WM-STGCN that leverages a Weighted adjacency matrix and Multi-scale temporal convolution in a spatio-temporal Graph Convolution Network to classify PD gaits from normal gaits and achieves up to 87.1% accuracy on 2D skeleton data. This method outperforms many existing methods including LSTM, KNN, DT, AdaBoost, and ST-GCN models [153]. [154] (ID = P-123) devises a static-dynamic temporal networks that consists of one-dimensional, two-dimensional convolutional networks, and attention mechanisms to classify PD patients based on foot force (vGRF time series) and achieves 96.7% accuracy.